OpenStudy (anonymous):

River C is 300 miles longer than River D. If the sum of their lengths is 5,600 miles, what is the length of each river?

OpenStudy (anonymous):

5,900

OpenStudy (anonymous):

your wrong @skate-4-life270

OpenStudy (anonymous):

i mean @sk8king14

OpenStudy (kamibug):

c = d + 300 c + d = 5600 d + 300 + d = 5600 2d + 300 = 5600 2d = 5300 d = 2650 <--- c + 2650 = 5600 c = 2950 <---

OpenStudy (anonymous):

oh sry I tryed

OpenStudy (tkhunny):

@hibba You are not understanding OpenStudy. The idea is for you to show YOUR work. Simply pasting the problem is strictly forbidden in the Code of Conduct. Have you considered solving correctly the system given to you by KamiBug? c = d + 300 c + d = 5600